An 15 g ice cube at -23˚C is put into a Thermos flask containing 150 cm3 of water at 17˚C. By how much has the entropy of the cube-water system changed when a final equilibrium state is reached? The specific heat of ice is 2200 J/kg K and that of liquid water is 4187 J/kg K. The heat of fusion of water is 333 × 103 J/kg.

Please give full explanation and the correct answer.
Answer #1

Mi = 11 g
Ti = -23 C = 250 K
Mw = 120 g
Tw = 17 C = 290 K
Hfus = 333 J/g.K
Cw = 4.187 J/g.K
Ci = 2.22 J/g.K

ΔEi + ΔEw = 0, ΔE = Ep+Ek+ΔU, Ep = Ek = 0
ΔUi + ΔUw = 0
ΔUi = -ΔUw
Mi*[Ci*(Tfus-Ti) + Hfus + Cw*(Tf-Tfus)] = Mw*C*(Tw-Tf)

Tf = 10676850-5556000/628050+62805 = 7.412 C

ΔS = ΔSi+ΔSw

ΔSi = Mi[(Ci*ln(Tfus/Ti) + Hfus/Tfus+ Cw ln(Tf/Tfus]

ΔSw = Mw*C ln(Tf/Tw)

ΔS =2.904+18.3+1.682-21.12 = 1.766 J/K